Hi again, How would I plot this?: y = -3/5 x + 1
Could you also explain how its done. The ny state test is comming and I still can't figure it out especially with fractions in it.
Thanks :)

Answer by ewatrrr(24785)


Hi,
Using the standard slope-intercept form for an equation of a line y = mx + b
where m is the slope and b the y-intercept.
y = -3/5 x + 1 is in this format with m = -(3/5) P(0,1) the y-intercept
m = -(3/5) tells us y decreases by 3 as x increases 5 or visa versa
Pt(0,1) the y-intercept tells us following are on the line with m = -(5/2):
Pt (5,-2)or as y decreases by 3, x inreases by 5
Pt (-5, 4) as y increases by 3, x decreases by 5